# NLP for Social Good: A Survey of Challenges, Opportunities, and Responsible Deployment

### v1

@@ 3. RQ3:
Inclusive NLP that accounts for demographic, linguistic, and accessibility diversity is key to reducing systemic inequalities. These efforts support SDG5 and SDG10, and address inequality—a top global concern ranked among the most severe short- and long-term risks (GR4).
Gender Bias
Gender bias in NLP—especially in hiring De-Arteaga et al. (2019), healthcare, and moderation—can reinforce stereotypes and harm marginalized groups. Initial work focused on static embeddings using gender direction Bolukbasi et al. (2016) and WEAT Caliskan et al. (2017), later extending to contextual tasks like coreference Zhao et al. (2018) and occupation classification De-Arteaga et al. (2019). Recent studies examine LLM-agent interactions Borah and Mihalcea (2024) to track how biases spread. Mitigation includes post-processing (e.g., fine-tuning, DPO), interpretability, causal inference Attanasio et al. (2023); Cai et al. (2024), and counterfactual or intervention-based prompting Plyler and Chi (2025). Bias is commonly assessed via fairness gaps (e.g., F1, error disparities) or associations (e.g., WEAT, Bias-Score).
Cultural and Linguistic Representation
NLP systems often fail to capture the full diversity of human cultures and languages. Culture’s complexity - including evolving norms, values, and worldviews - varies widely across communities (Saha et al., 2025). Despite culture’s crucial role in communication, AI models frequently reflect incomplete cultural perspectives, overrepresenting dominant or cultural linguistic narratives typically from high-resource languages or communities (Hershcovich et al., 2022a; Mihalcea et al., 2025).
Underrepresented Communities
Users with diverse accessibility needs remain underserved, continuing to face significant communication barriers Khanuja et al. (2023a). NLP has been used in a wide range of assistive applications, including as augmentative communication Park et al. (2022), text simplification Espinosa-Zaragoza et al. (2023), text-to-speech Kumar et al. (2023), speech recognition Li et al. (2022), braille processing Tejesh et al. (2025), image captioning and subtitling Stefanini et al. (2021), question answering Gurari et al. (2018), assistive chatbots Grassini et al. (2024), reading aids Wang et al. (2024d), and sign language translation Rust et al. (2024).
@@
Many issues revolve around binary gender assumptions and limited cross-linguistic/cultural generalizability Stanovsky et al. (2019); Adilazuarda (2024). Western-centric resources and lack of intersectional frameworks reinforce biases and marginalization  Kleinberg and Raghavan (2021); Sewunetie et al. (2024). Implicit biases and compounded discrimination, e.g., at the intersection of race, gender and disability, are underexplored Stanczak and Augenstein (2021); Guo and Caliskan (2021); Wald (2021). Data challenges persist due to scarce multimodal datasets, especially for Braille and sign language Hutchinson and Prabhakaran (2020); De Sisto et al. (2022). Limited interdisciplinary collaboration slows progress Kusters et al. (2020). Reliance on synthetic or prompt-generated data hampers real-world validity and hides systemic biases Venkit et al. (2025); Morales et al. (2024), while opaque data sourcing blocks accountability Bender and Friedman (2018).
@@
Inclusivity in NLP advances through participatory approaches centering marginalized voices, including co-designed fairness goals in gender bias mitigation Borah and Mihalcea (2024); Ma et al. (2023); Lauscher et al. (2022) and collaborative data collection with cultural experts and underrepresented groups Bird and Yibarbuk (2024); Newman-Griffis et al. (2024); UNICEF (2020); Hirmer et al. (2021). Solutions such as dynamic audit pipelines Park et al. (2023), lightweight model editing Park et al. (2023); Cai et al. (2024) and counterfactual data augmentation Zmigrod et al. (2019) help adapt models to sociocultural shifts. Fine-tuning LLMs can improve fairness and relevance Mai and Carson-Berndsen (2024); Bartl and Leavy (2024). Future work should consider identity compositionality Welch et al. (2020) and pluralistic alignment reflecting complex social affiliations Sorensen et al. (2024). Personalized, multimodal interaction design is vital for adaptive, accessible systems Paice et al. (2025); Wang et al. (2024d).
8 [Uncaptioned image]  Online Harms
@@
Recent social media platform statements [^5] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes highlight the limitations of AI-based content moderation at scale. Challenges include the subjective nature of moderation, regional regulations, and the culturally diverse and implicit nature of content (Ocampo et al., 2023). LLMs struggle with volatile topics without frequent fine-tuning (Roy et al., 2023), and efforts are hindered by unclear label taxonomies, biases in pre-trained models, and limited collaboration between lawmakers, platforms, and researchers (Yimam et al., 2024).
@@
Opportunities for future work include building low-resource, robust, and generalizable LLM moderation frameworks that are deployable in real time, requiring stronger collaboration between researchers, moderators, and policymakers (Munzert et al., 2025; Bui et al., 2025). A clearer, more widely accepted taxonomy of digital violence is needed to reduce ambiguity in label definitions (Yimam et al., 2024). The current focus on text-based hate overlooks multimodal content like video and audio, calling for LLMs with multimodal moderation capabilities. To address culturally diverse and implicit content, future systems should integrate explainability, fact-checking, and real-time tools such as RAGs and search APIs. Finally, moderation efforts must expand beyond hate speech to cover the full spectrum of digital violence, while also embracing more inclusive approaches that consider author and receiver demographics, and broader cultural and linguistic diversity.
@@
Ranked among the top five global risks in both short- and long-term scenarios (GR5), misinformation has a direct impact on the success of many of the SDGs. [^6] https://www.un.org/sites/un2.un.org/files/information-integrity-and-sdgs-en.pdf NLP approaches to misinformation include tasks such as fake news detection, rumor classification, stance detection, and fact checking (Oshikawa et al., 2020; Nakashole and Mitchell, 2014). Early methods used stylometric features, while modern systems use neural models, pre-trained LMs, and techniques like retrieval-augmented generation and prompt-based learning in multimodal and multilingual settings (Akhtar et al., 2023; Chen and Shu, 2024b).
@@
A key opportunity lies in developing human-centered evaluation methods tailored to real-world misinformation detection tasks (Das et al., 2023a). This would also benefit from interdisciplinary collaboration—for example, by integrating fundamental theories from social sciences and economics (Zhou and Zafarani, 2020), or by partnering with moderators, policymakers, and professional fact-checkers (Warren et al., 2025). Additionally, advancing the timeliness of ground truth through LLM-agents with access to web evidence and external knowledge bases can improve adaptability to fast-evolving misinformation. Developing domain-agnostic features that generalize across topics, languages, modalities, and time would help track shifting deceptive styles (Chen and Shu, 2024b) and identify check-worthy claims.
@@
AI harms are among the top global threats (GR5). NLP is part of the problem—but also part of the solution. We start with the black-box problem and follow the harm taxonomy by Weidinger et al. (2022) to explore NLP mitigation strategies.
The Black Box Problem
@@
Privacy, Safety and Malicous Uses
NLP approaches to harm reduction include learning to refuse or prevent memorization to protect personal data (Carlini et al., 2021; Liu et al., 2025b), adversarial training (Goyal et al., 2023), safe decoding (Xu et al., 2024b), authorship verification (Huang et al., 2024a), safety alignment (Bhardwaj et al., 2024), and red-teaming to expose vulnerabilities (Purpura et al., 2025).
Ungrounded Knowledge
@@
Socioeconomic and Environmental Harms
@@
Key barriers to prevent AI harms include transitioning from prototype to deployment which entails coping with data distributional shifts between “lab” and “field” data, and managing bias or subjective labels. Many current benchmarks often overlook low-resource or politically sensitive domains (Kim et al., 2025; Cho et al., 2025). Additionally, the lack of standardized practices for measuring real-world impact complicates evaluation beyond standard metrics. The effectiveness of the proposed tasks is often questioned  (Chen and Shu, 2024a; Kotonya and Toni, 2024), and there are very few rigorous and holistic evaluation frameworks Atanasova et al. (2023); Liang et al. (2023). There are also several governance problems and establishing equitable partnerships with nonprofits, where power imbalances or misaligned goals can hinder collaboration.
@@
Field trials of NLP systems under real-world conditions are necessary, especially in low-resource or use-case-specific settings, while encouraging participatory design and critical user engagement (Pataranutaporn et al., 2025). On the modeling side, advances in retrieval-augmented generation (RAG) and knowledge augmentation methods can improve reliability by enhancing credibility (Xu et al., 2024a; Chen et al., 2024, 2025). Multidisciplinary approaches such as logical relationships between inputs (Ayoobi et al., 2025; Freedman et al., 2025) and mechanistic interpretations of model behavior (Hou et al., 2023; Yu and Ananiadou, 2024) can further enhance transparency. New research directions could be implemented by studying LLM overreliance, manipulation and unfair distribution of benefits from model access. Embracing process-aware NLP enables alignment with domain logic and ethical goals through explainability and fairness-by-design (Bernardi et al., 2024; Zhuang et al., 2025). Finally, mapping existing frameworks for AI4SG and NLP4PI (Fairness, Accountability, and Transparency in Machine Learning (2020), FAT/ML; Floridi et al., 2021), suggested actions NIST (2024) and principles for trustworthy AI (OECD, 2024), can support the development of adaptable, unified guidelines for responsible NLP deployment.
@@ 11 Summary and Call to Action
To achieve the full potential of NLP4SG, the community should move beyond task-centric innovation toward impact-driven systems that are safe, inclusive, and globally relevant. This work reviewed nine key application areas, addressing RQ1 on existing impactful solutions with concrete examples of how modern NLP technologies have already contributed. At the same time, in addressing RQ2 on key open challenges, we find that areas like misinformation, online harms, and education have seen sustained attention in NLP, while domains such as poverty alleviation, environmental protection, peacebuilding are only starting to gain traction in response to real-world crises. In response to RQ3 concerning promising directions for future work, we outline a set of universal challenges, opportunities, and actionable recommendations to guide future research in NLP4SG:
Recurring Challenges: Across these diverse domains, we identified five systemic challenges that continue to hinder progress: (1) Data scarcity and representational bias, particularly affecting low-resource languages and marginalized groups; (2) Misaligned evaluation metrics that fail to capture human-centered qualities like empathy or cultural sensitivity; (3) Safety, privacy, and ethical concerns, which are magnified in high-stakes settings such as mental health or misinformation detection; (4) Limited generalization across cultures, modalities, and real-world conditions; and (5) Infrastructural gaps and institutional fragmentation, especially in underrepresented scenarios.
@@
Emerging Opportunities: Despite these barriers, we highlight several promising cross-cutting directions: (1) Multimodal and multilingual learning can help systems better reflect real-world diversity; (2) Human-AI collaboration enables more aligned, adaptive and interpretable NLP pipelines; (3) Participatory design and evaluation ensure that benchmarks and systems are co-developed with affected communities; (4) Retrieval-augmented and policy-aware methods offer new tools for building verifiable and context-sensitive applications; and (5) Explainability and AI literacy foster critical engagement and equitable access.
@@
Call to Action: To advance NLP4SG, we call on the community to: (1) develop joint benchmarks featuring multilingual, culturally diverse, and socially grounded data; (2) collaborate closely with domain experts, such as educators, health practitioners, and civil society organizations, to co-design evaluation frameworks that reflect end-user needs; (3) pursue human-centered methodologies instead of one-size-fits-all solutions. Progress depends on pluralistic, context-aware roadmaps that align with both local realities and global development goals. Finally, while modern LLMs offer significant potential, it is crucial to ensure their affordability and accessibility so they serve the public good rather than exacerbate existing inequalities.
@@
NLP has the tools to move beyond abstract benchmarking and toward socially responsive technologies designed with—and for—impacted communities. Realizing this vision requires not just technical innovation but also sustained interdisciplinary collaboration, inclusive practices, and a commitment to long-term global equity. We hope our findings can help earlier career researchers find their research niche and that more advanced researchers will have a fresh overview of the field to foster NLP4SG applications with a more interdisciplinary paradigm.
@@
This work is grounded in the belief that NLP research should be aligned with broader societal priorities and developed with care for its downstream impact. In proposing mappings between NLP directions, global goals, and risks, we are mindful of the potential for unintended consequences and overgeneralization. The open areas of work mentioned throughout the paper are just suggestive, and the practitioners, when working on them, should carefully evaluate the setting and downstream impacts.
We emphasize the importance of interdisciplinary collaboration and the inclusion of marginalized perspectives in shaping responsible NLP research. Where possible, we cite and build on existing initiatives in NLP for social good, and we support efforts to foreground equity, inclusivity, and transparency in both research framing and methodology. Finally, we acknowledge that AI assistants were used for proofreading this paper, a practice we consider ethically acceptable under appropriate usage.

### v2

@@ 1 Introduction
Inclusive and equitable language technologies are central to addressing systemic inequalities (disparities in how individuals or groups are represented, served, or affected), often reflecting socioeconomic, cultural, linguistic, and accessibility hierarchies. These efforts align with key sustainable development goals (SDG5, SDG10) and address inequality, ranked among the most severe societal risks in both the short and long term (GR4).
A growing body of survey work examines bias and fairness in NLP, cataloguing demographic bias (gender, race, ethnicity, age, sexual orientation, disability, and socioeconomic status) (Gupta et al., 2024), its origins (Brunet et al., 2019; Hovy and Prabhumoye, 2021), alongside detection, quantification and mitigation methods  Stanczak and Augenstein (2021); Devinney et al. (2022); Bartl et al. (2025); Gallegos et al. (2024). Studies have focused on bias in model representations Bolukbasi et al. (2016); Caliskan et al. (2017), creating bias detection benchmarks (Zhao et al., 2018; Parrish et al., 2022), and testing models in high-stakes settings De-Arteaga et al. (2019); Cross et al. (2024). More recent research addresses bias in agentic systems Borah and Mihalcea (2024), proposing mitigation strategies including post-processing, interpretability-driven and causal approaches Attanasio et al. (2023); Cai et al. (2024), as well as counterfactual prompting Plyler and Chi (2025).
Beyond demographic bias, inequalities also arise from the limited support of underrepresented communities with diverse accessibility needs Khanuja et al. (2023a). NLP has been applied to a wide range of assistive technologies, including augmentative communication Park et al. (2022), text simplification Espinosa-Zaragoza et al. (2023), text-to-speech and speech recognition Kumar et al. (2023); Li et al. (2022), braille processing Tejesh et al. (2025), image captioning and subtitling Stefanini et al. (2021), question answering Gurari et al. (2018), assistive chatbots Grassini et al. (2024), reading aids Wang et al. (2024d), and sign language translation Rust et al. (2024).
@@
Many challenges stem from simplifying assumptions, including binary gender representations, coarse socioeconomic proxies (Bassignana et al., 2025), and unequal access to digital resources Narayanan Venkit (2023). Limited cross-linguistic and cross-cultural generalizability further constrains model reliability Stanovsky et al. (2019); Adilazuarda (2024). Culture—encompassing evolving norms, values, and worldviews—varies widely across communities (Saha et al., 2025), yet NLP systems often reflect incomplete cultural perspectives, overrepresenting dominant linguistic narratives from high-resource languages and regions (Hershcovich et al., 2022a; Karamolegkou et al., 2024; Mihalcea et al., 2025). Western-centric resources and the lack of intersectional frameworks reinforce marginalization Kleinberg and Raghavan (2021); Sewunetie et al. (2024), while compounded forms of discrimination—arising at the intersection of race, gender, social status, and disability—remain underexplored Stanczak and Augenstein (2021); Guo and Caliskan (2021); Wald (2021). Progress is further hindered by scarce multimodal datasets, particularly for Braille and sign languages Hutchinson and Prabhakaran (2020); De Sisto et al. (2022); Karamolegkou et al. (2025b), limited interdisciplinary collaboration Kusters et al. (2020), and reliance on synthetic or prompt-generated data that obscures real-world biases Venkit et al. (2025); Morales et al. (2024). Finally, opaque data sourcing practices limit accountability and reproducibility Bender and Friedman (2018).
@@
Inclusivity in NLP advances through participatory approaches centering marginalized voices, including co-designed fairness goals in gender bias mitigation Borah and Mihalcea (2024); Ma et al. (2023); Lauscher et al. (2022) and collaborative data collection with cultural experts and underrepresented groups UNICEF (2020); Hirmer et al. (2021); Bird and Yibarbuk (2024); Newman-Griffis et al. (2024); Karamolegkou et al. (2024). Solutions such as dynamic audit pipelines Park et al. (2023), lightweight model editing Park et al. (2023); Cai et al. (2024) and counterfactual data augmentation Zmigrod et al. (2019) help adapt models to sociocultural shifts. Fine-tuning LLMs can improve fairness and relevance Mai and Carson-Berndsen (2024); Bartl and Leavy (2024). Future work should consider identity compositionality Welch et al. (2020) and pluralistic alignment reflecting complex social affiliations Sorensen et al. (2024). Personalized, multimodal interaction design is vital for adaptive, accessible systems Paice et al. (2025); Wang et al. (2024d).
8 [Uncaptioned image]  Digital Violence
@@
Recent social media platform statements [^5] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes highlight the limitations of AI-based content moderation. Challenges include the subjective nature of moderation, regional regulations, and the culturally diverse and implicit nature of content (Ocampo et al., 2023). LLMs struggle with volatile topics without frequent fine-tuning (Roy et al., 2023), and efforts are hindered by unclear label taxonomies, biases in pre-trained models, and limited collaboration between lawmakers, platforms, and researchers (Yimam et al., 2024).
@@
We argue that future work should build low-resource, robust, and generalizable LLM moderation frameworks deployable in real time, requiring stronger collaboration between researchers, moderators, and policymakers (Munzert et al., 2025; Bui et al., 2025). A shared taxonomy of digital violence is needed to reduce labeling ambiguity (Yimam et al., 2024). Current text-centric moderation overlooks multimodal content such as video and audio, highlighting the need for multimodal LLMs. To handle implicit and culturally diverse cases, future systems should integrate explainability, fact-checking, and real-time tools like RAGs and search APIs. Finally, moderation must move beyond hate speech to encompass the broader spectrum of digital violence—including harassment, bullying, spam, and abuse—through inclusive, context-aware systems that reflect diverse cultural, regional, and linguistic realities Moghaddam et al. (2025); Arora et al. (2023).
@@
Misinformation, the exposure to incorrect or misleading information, is ranked among the top five global risks in both short- and long-term scenarios (GR5), and has a direct impact on the success of many of the SDGs. [^6] https://www.un.org/sites/un2.un.org/files/information-integrity-and-sdgs-en.pdf NLP tools can be both part of the solution, but also part of the problem (see §10). There are many approaches to tackle misinformation with tasks such as fake news detection, rumor classification, stance detection, and fact checking (Oshikawa et al., 2020; Nakashole and Mitchell, 2014). Early methods used stylometric features, while modern systems use neural models, pre-trained LMs, and techniques like RAG and prompt-based learning in multimodal and multilingual settings (Akhtar et al., 2023; Chen and Shu, 2024b). Recently LLM agents are used for generating fact-checking articles (Sahnan et al., 2025), verifying complex claims (Chowdhury et al., 2025; Wang et al., 2025), and handling long-form or machine-generated texts (Xie et al., 2025; Boonsanong et al., 2025).
@@
A key opportunity lies in developing human-centered evaluation methods tailored to real-world tasks (Das et al., 2023a), and strengthened through interdisciplinary collaboration–integrating social and economic theories (Zhou and Zafarani, 2020) and partnerships with moderators, policymakers, and fact-checkers (Warren et al., 2025). Improving the timeliness of ground truth via LLM-agents that access web evidence and external knowledge bases can enhance adaptability to fast-evolving content. Developing domain-agnostic features generalizable across topics, languages, and modalities can help detect shifting deceptive styles (Chen and Shu, 2024b) and identify check-worthy claims. The growing threat of LLM-generated misinformation remains a critical frontier (Liu et al., 2025a; Gabriel et al., 2024), underscoring the need for adaptive, accountable, and transparent detection systems.
@@
AI harms, referring to negative impacts stemming from the design, deployment, or use of AI systems, are among the top global threats (GR5). NLP is part of the problem—but also part of the solution. We first consider the black-box nature of AI systems and then follow the harm taxonomy by Weidinger et al. (2022) to structure our discussion and explore NLP-based mitigation strategies.
The Black Box Problem.
@@
Privacy, Safety and Malicous Uses.
NLP approaches to harm reduction include learning to refuse or prevent memorization to protect personal data (Carlini et al., 2021; Liu et al., 2025c), adversarial training (Goyal et al., 2023), safe decoding (Xu et al., 2024b), authorship verification (Huang et al., 2024a), safety alignment (Bhardwaj et al., 2024), and red-teaming to expose vulnerabilities (Purpura et al., 2025).
Ungrounded Knowledge.
@@
Socioeconomic and Environmental Harms.
@@
Key barriers to prevent AI harms include transitioning from prototype to deployment which entails coping with data distributional shifts between “lab” and “field” data, and managing bias or subjective labels. Many current benchmarks often overlook low-resource or politically sensitive domains (Kim et al., 2025; Cho et al., 2025). Additionally, the lack of standardized practices for measuring real-world impact complicates evaluation beyond standard metrics. The effectiveness of the proposed tasks is often questioned  (Chen and Shu, 2024a; Kotonya and Toni, 2024), and there are very few rigorous and holistic evaluation frameworks Atanasova et al. (2023); Liang et al. (2023). There are also several governance problems and establishing equitable partnerships with nonprofits, where power imbalances or misaligned goals can hinder collaboration. Lastly, while NLP methods can mitigate certain AI harms, many of them—such as overreliance or erosion of human agency—cannot be resolved through technical solutions alone.
@@
Field trials of NLP systems under real-world conditions are necessary, especially in low-resource or use-case-specific settings, while encouraging participatory design and critical user engagement (Pataranutaporn et al., 2025). On the modeling side, advances in retrieval-augmented generation (RAG) and knowledge augmentation methods can improve reliability by enhancing credibility (Xu et al., 2024a; Chen et al., 2024, 2025). Multidisciplinary approaches such as logical relationships between inputs (Ayoobi et al., 2025; Freedman et al., 2025) and mechanistic interpretations of model behavior (Hou et al., 2023; Yu and Ananiadou, 2024) can further enhance transparency. Additional studies could examine LLM overreliance, manipulation, the uneven distribution of benefits from model access, as well as tracing gaps and overlaps between professional and laypeople AI concerns (Karamolegkou et al., 2025a). Embracing process-aware NLP enables alignment with domain logic and ethical goals through explainability and fairness-by-design (Bernardi et al., 2024; Zhuang et al., 2025). Finally, mapping existing frameworks for AI4SG and NLP4PI (Fairness, Accountability, and Transparency in Machine Learning (2020), FAT/ML; Floridi et al., 2021), suggested actions NIST (2024) and principles for trustworthy AI (OECD, 2024), can support the development of adaptable, unified guidelines for responsible NLP deployment.
@@ 11 Summary and Call to Action
To achieve the full potential of NLP4SG, the community should move beyond task-centric innovation toward impact-driven systems that are safe, inclusive, and globally relevant. This work reviewed nine key application areas, addressing RQ1 on existing impactful solutions with concrete examples of how modern NLP technologies have already contributed. Complementing this review, our quantitative analysis of 47K ACL papers showed that research on AI harms and inclusion has grown most rapidly, while domains like poverty, peacebuilding, and environment remain underexplored (Figure 1 and Appendix A). At the same time, in addressing RQ2 on key open challenges, we find that areas like misinformation, online harms, and education have seen sustained attention in NLP, while domains such as poverty alleviation, environmental protection, and peacebuilding are only starting to gain traction in response to real-world crises. Addressing RQ3 concerning promising directions for future work, we outline a set of universal challenges, opportunities, and actionable recommendations to guide future research in NLP4SG:
@@
Emerging Challenges and Opportunities: Despite barriers such as data scarcity and representational bias, misaligned evaluation metrics, and persistent safety, privacy, and ethical concerns, alongside ongoing infrastructural gaps, we identify several promising directions for progress. (1) Multilingual and multicultural learning can help systems better reflect real-world diversity including not only mainstream languages, but caring about local language communities equally; (2) Human–AI collaboration enables more adaptive and interpretable NLP pipelines; (3) Participatory design and evaluation ensure that systems are co-developed with affected communities; (4) Retrieval-augmented and policy-aware methods provide tools for verifiable, context-sensitive applications; and (5) Explainability and AI literacy foster critical engagement and equitable access.
@@
Call to Action: To advance NLP4SG, we call on the community to: (1) Develop joint benchmarks featuring multilingual, culturally diverse, and socially grounded data; (2) Collaborate closely with domain experts, such as educators, health practitioners, and civil society organizations, to co-design evaluation frameworks that reflect end-user needs. (3) Pursue human-centered methodologies instead of one-size-fits-all solutions. Progress depends on pluralistic, context-aware roadmaps that align with both local realities and global development goals. (4) Finally, while modern LLMs offer significant potential, it is crucial to ensure their affordability and accessibility so they serve the public good rather than exacerbate existing inequalities. Assess whether deploying a large, one-size-fits-all generalist LLM is truly necessary, or whether more efficient, and more environmentally sustainable NLP solutions would be preferable.
@@
NLP has the tools to move beyond abstract benchmarking and toward socially responsive technologies designed with—and for—impacted communities. Realizing this vision requires not just technical innovation but also sustained interdisciplinary collaboration, inclusive practices, and a commitment to long-term global equity. We hope our findings can help researchers early in their careers to find their research niche and that more advanced researchers will have a fresh overview of the field to foster NLP4SG applications with a more interdisciplinary paradigm.
@@
This work is grounded in the belief that NLP research should be aligned with broader societal priorities and developed with care for its downstream impact. All ACL materials used for our statistical analysis are licensed using the Creative Commons 3.0 BY-NC-SA. In proposing mappings between NLP directions, global goals, and risks, we are mindful of the potential for unintended consequences. The open areas of work mentioned throughout the paper are just suggestive, and the practitioners, when working on them, should carefully evaluate the setting and downstream impacts.
We emphasize the importance of interdisciplinary collaboration and the inclusion of marginalized perspectives in shaping responsible NLP research. Where possible, we cite and build on existing initiatives in NLP for social good, and we support efforts to foreground equity, inclusivity, and transparency in both research framing and methodology. Throughout this paper, we emphasize that NLP systems should augment rather than replace human expertise, particularly in sensitive or high-stakes contexts. Finally, we acknowledge that AI assistants were used for proofreading this paper, a practice we consider ethically acceptable under appropriate usage.